# Enhancing Fairness in Unsupervised Graph Anomaly Detection through Disentanglement

### v1

@@ Abstract.
Graph anomaly detection (GAD) is becoming increasingly crucial in various applications, ranging from financial fraud detection to fake news detection. However, current GAD methods largely overlook the issue of fairness, which might result in discriminatory decisions skewed toward certain demographic groups defined on sensitive attributes (e.g., gender, religion, ethnicity, etc.). This greatly limits the applicability of these methods in real-world scenarios in light of societal and ethical restrictions. To address this critical gap, we make the first attempt to integrate fairness with utility in GAD decision-making. Specifically, we devise a novel DisEntangle-based FairnEss-aware aNomaly Detection framework on the attributed graph, named DEFEND. DEFEND first introduces disentanglement in GNNs to capture informative yet sensitive-irrelevant node representations, effectively reducing societal bias inherent in graph representation learning. Besides, to alleviate discriminatory bias in evaluating anomalous nodes, DEFEND adopts a reconstruction-based anomaly detection, which concentrates solely on node attributes without incorporating any graph structure. Additionally, given the inherent association between input and sensitive attributes, DEFEND constrains the correlation between the reconstruction error and the predicted sensitive attributes. Our empirical evaluations on real-world datasets reveal that DEFEND performs effectively in GAD and significantly enhances fairness compared to state-of-the-art baselines. To foster reproducibility, our code is available at https://github.com/AhaChang/DEFEND.
Graph Anomaly Detection, Fairness, Graph Neural Networks
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXisbn: 978-1-4503-XXXX-X/18/06ccs: Mathematics of computing Graph algorithmsccs: Security and privacy Software and application security
1. Introduction
@@
Graph Anomaly Detection (GAD), which aims to identify nodes that deviate significantly from the majority of nodes, has attracted wide attention in various domains, including fraudster detection in financial networks (Zhang et al., 2022; Huang et al., 2022) and spammer detection in social networks (Li et al., 2019; Wu et al., 2020). The advancement of Graph Neural Networks (GNNs) (Kipf and Welling, 2016; Hamilton et al., 2017; Veličković et al., 2017) has significantly propelled the development of GNN-based GAD methods (Ding et al., 2019; Chai et al., 2022; Kim et al., 2023; He et al., 2023), enhancing their ability to accurately identify anomalies on graphs.
@@
Given the wide-ranging applications of graph anomaly detection, particularly within high-stakes domains, its fairness cannot be overlooked. Specifically, unfair decisions that skew toward certain demographic groups associated with sensitive attributes (e.g., gender, religion, and ethnicity) might cause profound societal and ethical concerns. For example, in the realm of social networks (e.g., Reddit and Twitter), anomalous users (e.g., spreading misinformation or engaging in fake account interactions) might be subjected to strict investigation and even permanent account suspension. In such scenarios, biased decisions could result in disproportionately targeting certain groups while inadvertently neglecting others. This undermines the effectiveness and reliability of anomaly detection systems and raises critical ethical concerns. To address the issues above, several fairness-aware anomaly detection methods have been proposed (Deepak and Abraham, 2020; Song et al., 2021; Zhang and Davidson, 2021; Shekhar et al., 2021), aiming to balance fairness and utility in anomaly detection. However, these methods predominantly focus on independent and identically distributed (i.i.d.) data, failing to account for the societal bias in graph data.
@@
Societal bias within graph data poses a significant challenge to achieving fairness in graph learning tasks (Dai and Wang, 2021; Zhu et al., 2023). The bias manifests not only in node attributes but also in the topological structure of the graph. From one aspect, sensitive attributes are likely inherently spread across other attributes (e.g., the geographic location might be correlated with religion and ethnicity) (Deepak and Abraham, 2020). Moreover, since nodes with similar attributes are more likely to form connections, the graph topology is actually influenced by sensitive attributes (Rahman et al., 2019; Spinelli et al., 2021). Additionally, recent studies (Dai and Wang, 2021; Wang et al., 2022; Zhu et al., 2023) have demonstrated that the biased topology coupled with the message-passing mechanism in GNNs may inherit or even amplify the inherent societal bias in graphs. More specifically, aggregating features from neighboring nodes that share identical sensitive attributes may result in representations that amplify the features of the demographic group. Consequently, such homogeneous representations can potentially affect the fairness of the decision-making process in GNNs.
@@
Over the past years, many existing works have explored fairness in graphs for GNN-based methods (Li et al., 2021; Ma et al., 2022; Kim et al., 2022; Song et al., 2022). A common strategy involves eliminating sensitive information from the training graph and implementing the debiased graph for the target tasks (Spinelli et al., 2021; Rahman et al., 2019; Dong et al., 2022). In this manner, the debiased graph is expected to retain less potentially sensitive information while preserving all non-sensitive characteristics. However, the efficacy of such debiasing techniques in GAD tasks is not yet fully established. A primary concern is the potential overlap between characteristics of anomalies associated with nonsensitive attributes and those of demographic groups linked to sensitive attributes. For instance, an edge might indicate that nodes share the same sensitive attributes while displaying anomalous behavior. This overlap poses a challenge in concurrently mitigating bias and preserving the integrity of anomalies. Another prevalent strategy involves training fair GNNs to perform the target tasks in a way that is independent of sensitive attributes (Dai and Wang, 2021; Zhu et al., 2023). These methods are dedicated to end-to-end supervised node classification, where target labels are available during training. Therefore, directly applying them to unsupervised graph anomaly detection is challenging, primarily due to the absence of ground truth labels that indicate whether a node is anomalous.
As an alternative, fair representation learning has been proposed to learn representations free from sensitive attributes while maintaining essential information for decision-making in downstream tasks (Zemel et al., 2013; Liu et al., 2022b). Disentangled representation learning provides a novel perspective in fair representation learning, enabling the simultaneous maintenance of sensitive-relevant and irrelevant information, which are separated into independent subspaces. While disentangled fair representation learning has shown promise in benefiting fairness in image classification tasks (Creager et al., 2019; Kim et al., 2021; Oh et al., 2022), its applicability on unsupervised graph anomaly detection is non-trivial. Different from the i.i.d. image data, graph data contains complex interrelations between nodes, which are represented in the graph topology. The first challenge involves effectively encoding both the attributes of nodes and the topology of the graph into a disentangled i.i.d. node representation while preserving the information essential for anomaly detection. Furthermore, an additional challenge lies in achieving fair anomaly detection with obtained fair representation. In prior studies, disentangled fair representation is used for supervised downstream tasks. However, we focus on unsupervised anomaly detection, which relies on reconstruction error to make the decision. Considering the correlation between sensitive attributes and other attributes, as well as the connections influenced by sensitive attributes, direct reconstruction of the original graph may result in biased decisions.
@@
In this paper, we propose a novel DisEntangle-based FairnEss-aware aNomaly Detection framework on the attributed graph, named DEFEND. We first introduce disentangled fair representation learning on graphs to capture node representations that are both informative and independent of sensitive attributes. Specifically, the disentangled graph encoder can effectively separate sensitive-relevant and -irrelevant node representations into independent subspaces with the guidance of a learnable adversary. To alleviate discriminatory bias in detecting anomalous nodes, we deploy an additional decoder that reconstructs node attributes from the well-trained disentangled encoder. Given the inherent bias in graph topology, anomaly detection relies solely on node attributes. Furthermore, to mitigate the effect of the interrelation between raw node attributes and sensitive attributes, we introduce a fairness constraint, ensuring the reconstruction error remains invariant to the predicted sensitive attributes.
In summary, the main contributions of this paper are as follows:
@@
  • DEFEND employs constrained reconstruction error coupled with a disentangled encoder for fair anomaly detection.
@@
Figure 1. An overview of proposed DEFEND framework. (Left) Disentangled fair representation learning. The disentangled graph encoder $f_{e}$ can separate sensitive-irrelevant representations $\mathbf{Z}_{x}$ and sensitive-relevant representations $\mathbf{Z}_{s}$ in latent space. (Right) Reconstruct-based graph anomaly detection. The constrained reconstruction error between $\mathbf{X}$ and $\tilde{\mathbf{X}}$ are used to identify anomalies. ❄ means fixing model parameters.
2. Preliminaries
2.1. Problem Definition
Let $\mathcal{G}=(\mathcal{V},\mathbf{A},\mathbf{X},\mathbf{S})$ be an attributed graph with $N$ nodes and $E$ edges, where $\mathcal{V}=\{v_{1},\dots,v_{N}\}$ is the set of nodes. The adjacency matrix is denoted as $\mathbf{A}\in{\{0,1\}}^{N\times N}$, where $\mathbf{A}_{ij}=1$ if there exists an edge between $v_{i}$ and $v_{j}$, otherwise, $\mathbf{A}_{ij}=0$. $\mathbf{X}\in\mathbb{R}^{N\times d}$ represents the node attribute matrix, while $\mathbf{S}\in\mathbb{R}^{N\times m}$ represents the sensitive attribute matrix (e.g., age, gender, religion, ethnicity, etc.). Here, $m$ denotes the total number of sensitive attributes. To simplify the problem, in this work, we mainly focus on a single binary sensitive attribute, denoted as $\mathbf{S}\in{\{0,1\}}^{N\times 1}$. We can easily extend our method to more complicated settings (Creager et al., 2019; Deepak and Abraham, 2020). We discuss more details in Section 3.5. Node $v_{i}$ and $v_{j}$ belong to the same demographic group if $s_{i}=s_{j}$. The goal of fair graph anomaly detection is to provide unbiased prediction against sensitive attributes while achieving satisfactory accuracy simultaneously.
2.2. Fairness Metrics
Following (Agarwal et al., 2021; Wang et al., 2022; Zhu et al., 2023), we utilize two extensively used metrics to evaluate the fairness of the model among demographic groups, including Demographic Parity (Dwork et al., 2012) and Equal Opportunity (Hardt et al., 2016). Notably, a lower value in Demographic Parity and Equal Opportunity reflects a higher level of fairness achieved by the model.
2.2.1. Demographic parity.
Demographic parity (DP) dictates the equal predicted probability across demographic groups. In GAD tasks, it indicates a consistent rate of anomaly detection across all demographic groups. It can be formulated as:
(1) $\Delta_{DP}=|P(\hat{y}=1|s=0)-P(\hat{y}=1|s=1)|,$
@@ 2.2. Fairness Metrics
2.2.2. Equal Opportunity.
Equal Opportunity (EO) requires the same true positive rates of identifying anomalies for each demographic group. Specifically, it implies that each demographic group should have an equivalent probability of correctly identifying true anomalies. It can be defined as:
(2) $\Delta_{EO}=|P(\hat{y}=1|s=0,y=1)-P(\hat{y}=1|s=1,y=1)|,$
where $y\in\{0,1\}$ is the ground truth anomaly label, and $y_{i}=1$ denotes node $v_{i}$ is a true anomaly.
3. Proposed Method
In this section, we describe the proposed DEFEND in detail. Initially, we provide an overview of DEFEND in Section 3.1. Then, we describe the disentangled representation learning phase in Section 3.2, and the anomaly detection phase in Section 3.3. Next, the training and inference processes of DEFEND are elaborated upon, alongside an analysis of its computational complexity in Section 3.4. Finally, we explore the generalization capabilities of DEFEND across diverse sensitive attributes in Section 3.5.
3.1. Overview
@@
We provide an overview of DEFEND, which aims to identify anomalies without skewed towards any demographic groups defined on the given sensitive attributes. As depicted in Figure 1, there are two major phases in DEFEND. Firstly, as shown in the left part of Figure 1, the fair disentangled representation learning phase separates sensitive-relevant and sensitive-irrelevant representations in the latent space. Different from the i.i.d. data, nodes in a graph are connected by edges. To effectively encode these dependencies, we propose to learn the node representations by a variational graph autoencoder. Specifically, a GNN encoder $f_{e}$ maps the node attribute $\mathbf{X}$ and graph topology $\mathbf{A}$ of the input graph $\mathcal{G}$ into independent subspaces, i.e., sensitive-relevant and sensitive-irrelevant subspaces. Subsequently, two decoders $f_{a}$ and $f_{x}$ are employed to generate the adjacency matrix $\hat{\mathbf{A}}$ and the node attributes $\hat{\mathbf{X}}$, respectively. This reconstruction enables node representations to further learn the topological dependencies from self-supervision signals. To promote the independence of sensitive-irrelevant representations $\mathbf{Z}_{x}$ from sensitive-relevant representations $\mathbf{Z}_{s}$, we include a learnable adversary $g_{\omega}$. Besides minimizing reconstruction error to obtain informative representations, we also endeavor to accurately infer sensitive attributes $\mathbf{S}$ from $\mathbf{Z}_{s}$. Ideally, $\mathbf{Z}_{x}$ contains nearly no sensitive information with a well-trained $f_{e}$. Secondly, the unsupervised graph anomaly detection phase aims to identify anomalies according to the sensitive-irrelevant representations, as shown in the right part of Figure 1. However, the reconstruction-based anomaly detection can incur bias from the reconstruction target. As such, the decoder $f_{\phi}$ only reconstructs node attributes from sensitive-irrelevant representations $\mathbf{Z}_{x}$ through Multi-Layer Perceptron (MLP) without the involvement of graph topology. Moreover, as node attributes $\mathbf{X}$ inherently have potential correlations with sensitive attributes $\mathbf{S}$, we propose to further constrain the correlation in the loss function.
3.2. Disentangled Fair Representation Learning
@@
Due to the potential bias existing in node attributes and graph topology, GNNs might amplify sensitive information when generating node representations with the message-passing process (Dai and Wang, 2021; Zhu et al., 2023). Informative sensitive-irrelevant representations are essential for fair and accurate decision-making in downstream tasks, which thoroughly considers all information pertinent to each node in terms of node and structure levels. Previous theoretical insights and empirical evidence have highlighted the effectiveness of disentanglement in generating sensitive-irrelevant representations and augmenting fairness in downstream tasks like image classification (Creager et al., 2019; Oh et al., 2022). We posit that disentangling sensitive and non-sensitive representations for graph data is also feasible to obtain informative yet sensitive-irrelevant node representations, which can be used for graph anomaly detection. The details of disentangled representation learning in DEFEND are described as follows.
3.2.1. Disentangled Graph Encoder.

### v2

@@ Abstract
Graph anomaly detection (GAD) is becoming increasingly crucial in various applications, ranging from financial fraud detection to fake news detection. However, current GAD methods largely overlook the fairness problem, which might result in discriminatory decisions skewed toward certain demographic groups defined on sensitive attributes (e.g., gender). This greatly limits the applicability of these methods in real-world scenarios in light of societal and ethical restrictions. To address this critical gap, we make the first attempt to integrate fairness with utility in GAD decision-making. Specifically, we devise a novel DisEntangle-based FairnEss-aware aNomaly Detection framework on the attributed graph, named DEFEND. DEFEND first introduces disentanglement in GNNs to capture informative yet sensitive-irrelevant node representations, effectively reducing bias inherent in graph representation learning. Besides, to alleviate discriminatory bias in evaluating anomalies, DEFEND adopts a reconstruction-based method, which concentrates solely on node attributes and avoids incorporating biased graph topology. Additionally, given the inherent association between sensitive-relevant and -irrelevant attributes, DEFEND further constrains the correlation between the reconstruction error and predicted sensitive attributes. Empirical evaluations on real-world datasets reveal that DEFEND performs effectively in GAD and significantly enhances fairness compared to state-of-the-art baselines. Our code is available at https://github.com/AhaChang/DEFEND.
1 Introduction
@@
Graph Anomaly Detection (GAD), which aims to identify nodes that deviate significantly from the majority of nodes, has attracted wide attention in various domains, including fraudster detection in financial networks (Zhang et al., 2022; Huang et al., 2022) and spammer detection in social networks (Li et al., 2019; Wu et al., 2020). The advancement of Graph Neural Networks (GNNs) (Kipf & Welling, 2016; Hamilton et al., 2017; Veličković et al., 2017) has significantly enhanced the ability of GNN-based GAD methods to accurately identify anomalies (Ding et al., 2019; Chai et al., 2022; Kim et al., 2023; He et al., 2023). However, a recent study (Neo et al., 2024) reveal a concerning trend: current GAD methods exhibit substantial bias in decision.
@@
Given the wide-ranging applications of GAD, particularly within high-stakes domains, the fairness problem cannot be overlooked. Unfair decisions that skew toward certain demographic groups associated with sensitive attributes (e.g., gender, religion, ethnicity, etc.) might cause profound societal and ethical concerns. For example, in the realm of social networks (e.g., Reddit and Twitter), anomalous users (e.g., spreading misinformation or engaging in fake account interactions) might undergo strict investigation and even permanent account suspension. In such scenarios, biased decisions could result in unfairly focusing on certain groups while inadvertently neglecting others. This undermines the effectiveness and reliability of anomaly detection systems and raises critical ethical concerns. To balance fairness and utility in anomaly detection, several methods have been proposed (Deepak & Abraham, 2020; Song et al., 2021; Zhang & Davidson, 2021; Shekhar et al., 2021). These methods strive to optimize the balance between fairness and anomaly detection performance in the absence of ground truth labels, which presents a fundamental challenge in unsupervised anomaly detection. Nevertheless, they primarily focus on independent and identically distributed data, thereby overlooking the societal bias in graphs, which manifests in both node attributes and graph topology.
@@
The bias in graphs poses a significant challenge to achieving fairness in graph-related tasks (Dai & Wang, 2021; Zhu et al., 2023). First, sensitive attributes are inherently spread across other attributes (Deepak & Abraham, 2020; Oh et al., 2022), so directly removing them is insufficient to ensure fairness (Neo et al., 2024). For example, the geographic location might correlate with religion and ethnicity. Second, since nodes with similar attributes are more likely to form connections, graph topology is also influenced by sensitive attributes (Rahman et al., 2019; Spinelli et al., 2021). Third, biased topology coupled with the message-passing mechanism in GNNs may inherit or even amplify the inherent bias in graphs (Dai & Wang, 2021; Wang et al., 2022; Zhu et al., 2023). Specifically, representations aggregated from neighboring nodes that share identical sensitive attributes may amplify features of the demographic group, thereby potentially affecting the fairness of the decision-making in GNNs.
@@
Many efforts have been made to explore fairness for GNN-based methods (Li et al., 2021; Ma et al., 2022; Kim et al., 2022; Song et al., 2022). A common strategy involves eliminating sensitive information from the training graph and implementing the debiased graph for target tasks (Spinelli et al., 2021; Rahman et al., 2019; Dong et al., 2022). However, it is non-trivial to concurrently mitigate bias and preserve the integrity of anomalies, considering the overlap between the features of anomalies unrelated to sensitive attributes and the features of demographic groups linked to sensitive attributes. For instance, an edge might indicate that two nodes share the same sensitive attributes while displaying an anomalous connection. Another prevalent strategy involves training fair GNNs to perform the target task independently of sensitive attributes (Dai & Wang, 2021; Zhu et al., 2023). These methods are dedicated to end-to-end supervised node classification, where ground-truth labels are available. Thus, applying them to unsupervised GAD is challenging due to the absence of labels for anomaly detection.
@@
In this paper, we propose a novel DisEntangle-based FairnEss-aware aNomaly Detection framework on attributed graphs, named DEFEND. In the first stage, to address the societal bias embedded in both node attributes and graph topology, we introduce disentangled fair representation learning on graphs to capture node representations that are both informative and independent of sensitive attributes. Specifically, the disentangled graph encoder can effectively separate sensitive-relevant and sensitive-irrelevant representations into independent subspaces with the guidance of a learnable adversary. In the second stage, given the absence of ground truth labels and the inherent complex bias in graph topology, we implement an additional decoder that reconstructs node attributes from the well-trained disentangled encoder, utilizing the reconstruction error as the anomaly score. To further alleviate discriminatory bias in detecting anomalies, we introduce a fairness constraint that enforces invariance of reconstruction errors across different sensitive attribute groups, effectively mitigating the influence of correlations between sensitive and the rest attributes. Our main contributions are summarized as follows:
@@
  • DEFEND employs constrained reconstruction error coupled with a disentangled graph encoder for fair anomaly detection on graphs.
@@
2 Preliminaries
2.1 Problem Definition
Let $\mathcal{G}=(\mathcal{V},\mathbf{A},\mathbf{X},\mathbf{S})$ be an attributed graph with $N$ nodes and $E$ edges, where $\mathcal{V}=\{v_{1},\dots,v_{N}\}$ is the set of nodes. $\mathbf{y}\in{\{0,1\}}^{N}$ denotes the anomaly labels, where 1 indicates an anomalous node, and $\hat{\mathbf{y}}$ denotes the predicted labels. The adjacency matrix is denoted as $\mathbf{A}\in{\{0,1\}}^{N\times N}$, where $\mathbf{A}_{ij}=1$ if there exists an edge between $v_{i}$ and $v_{j}$, otherwise, $\mathbf{A}_{ij}=0$. $\mathbf{X}\in\mathbb{R}^{N\times d}$ represents the observed node attribute matrix, while $\mathbf{S}\in\mathbb{R}^{N\times m}$ represents the sensitive attribute matrix (e.g., gender, religion, ethnicity, etc.). $v_{i}$ and $v_{j}$ belong to the same demographic group if $s_{i}=s_{j}$. Here, $m$ is the total number of sensitive attributes. As described in (Deepak & Abraham, 2020; Sarhan et al., 2020; Oh et al., 2022), the sensitive attribute $\mathbf{S}$ is correlated with both the observed attribute $\mathbf{X}$ and labels $\mathbf{y}$ on many real-world datasets. The goal of fair graph anomaly detection is to provide unbiased prediction against sensitive attributes while achieving satisfactory accuracy simultaneously. To simplify the problem, in this work, we mainly focus on a single binary sensitive attribute, i.e., $\mathbf{S}\in{\{0,1\}}^{N\times 1}$. We can easily extend our method to more complicated settings as previous studies (Creager et al., 2019; Deepak & Abraham, 2020). More details are discussed in Section 3.3.
2.2 Fairness Metrics
Following (Agarwal et al., 2021; Wang et al., 2022; Zhu et al., 2023), we utilize two widely used metrics to evaluate the fairness of models among demographic groups. Demographic Parity Dwork et al. (2012) dictates the equal predicted probability across demographic groups. It ensures predictions are statistically unbiased to sensitive attributes, e.g., if gender is a sensitive attribute, $\Delta_{DP}$ implies that the probability of individuals from different genders being classified as anomalous should be identical.
$\Delta_{DP}=|P(\hat{y}=1|s=0)-P(\hat{y}=1|s=1)|,$ (1)
@@ 2.2 Fairness Metrics
Equal Opportunity (Hardt et al., 2016) requires the same true positive rates of identifying anomalies for each demographic group. Considering gender as a sensitive attribute, $\Delta_{EO}$ encourages individuals from different genders to have an equal probability of being correctly identified as anomalous. Given the ground-truth label $y\in\{0,1\}$ where $y_{i}=1$ denotes $v_{i}$ is a true anomaly, $\Delta_{EO}$ can be defined as:
$\Delta_{EO}=|P(\hat{y}=1|s=0,y=1)-P(\hat{y}=1|s=1,y=1)|.$ (2)
3 Proposed Method
We now introduce DEFEND, which aims to identify anomalies without skew towards demographic groups defined on sensitive attributes. The disentangled representation learning separates sensitive-relevant and sensitive-irrelevant representations in the latent space (see Section 3.1). Fair anomaly detection is achieved through constrained reconstruction error using the disentangled representations (see Section 3.2). We then discuss its generalization capability across diverse sensitive attributes in Section 3.3. We introduce the training and inference processes in Appendix A and provide a computational complexity analysis in Appendix B.
@@ 3 Proposed Method
Figure 1 illustrates the overall workflow of DEFEND, which contains two major phases. Firstly, as shown in the left part of Figure 1, the disentangled fair representation learning phase separates sensitive-relevant representations $\mathbf{Z}_{x}$ and sensitive-irrelevant representations $\mathbf{Z}_{s}$ in the latent space. Specifically, a disentangled graph encoder $f_{e}$ maps the node attribute $\mathbf{X}$ and graph topology $\mathbf{A}$ into independent sensitive-relevant and -irrelevant subspaces. Subsequently, $f_{a}$ and $f_{x}$ decode adjacency matrix $\hat{\mathbf{A}}$ and node attributes $\hat{\mathbf{X}}$, respectively. Ideally, $\mathbf{Z}_{x}$ contains no sensitive information with a well-trained $f_{e}$. To promote the independence of $\mathbf{Z}_{x}$ from $\mathbf{Z}_{s}$, we include a learnable adversary $g_{\omega}$. Besides minimizing reconstruction error to obtain informative representations, we also endeavor to accurately infer sensitive attributes $\mathbf{S}$ from $\mathbf{Z}_{s}$. As shown in the right part of Figure 1, the unsupervised graph anomaly detection phase identifies anomalies based on sensitive-irrelevant representations $\mathbf{Z}_{x}$. However, the reconstruction of attributes and structures may lead to biased outcomes in decision-making, considering biases inherent in both attributes and structures. As such, the decoder $f_{\phi}$ only reconstructs node attributes from $\mathbf{Z}_{x}$ based on Multi-Layer Perceptron (MLP) without involving graph topology. Moreover, as node attributes $\mathbf{X}$ inherently have potential correlations with sensitive attributes $\mathbf{S}$, we further constrain the correlation between the reconstruction error and the predicted sensitive attributes.
Figure 1: An overview of proposed DEFEND framework. (Left) Disentangled fair representation learning. The disentangled graph encoder $f_{e}$ can separate sensitive-irrelevant representations $\mathbf{Z}_{x}$ and sensitive-relevant representations $\mathbf{Z}_{s}$ in latent space. (Right) Reconstruct-based graph anomaly detection. The constrained reconstruction error between $\mathbf{X}$ and $\tilde{\mathbf{X}}$ are used to identify anomalies. ❄ means fixing model parameters.
3.1 Disentangled Fair Representation Learning
@@
Informative sensitive-irrelevant representations are crucial for fair and accurate decision-making in GAD, which thoroughly considers features of nodes at both node and structural levels. However, since the potential bias in node attributes and graph topology, GNNs might amplify sensitive information when generating node representations with the message-passing mechanism (Dai & Wang, 2021; Zhu et al., 2023). Previous theoretical insights and empirical evidence have highlighted the effectiveness of disentangled representation learning in separating sensitive-irrelevant representations for augmenting fairness in downstream tasks like image classification (Creager et al., 2019; Oh et al., 2022). We posit that disentangled representation learning is also feasible to provide informative yet sensitive-irrelevant node representations for GAD. The details of disentangled fair representation learning on graph-structured data in DEFEND are described as follows.
3.1.1 Disentangled Graph Encoder